# ExeCoder: Empowering Large Language Models with Executability Representation for Code Translation

### v2

@@
Abstract.
Code translation is a crucial activity in the software development and maintenance process, and researchers have recently begun to focus on using pre-trained large language models (LLMs) for code translation. However, existing LLMs only learn the contextual semantics of code during pre-training, neglecting executability information closely related to the execution state of the code, which results in unguaranteed code executability and unreliable automated code translation. To address this issue, we propose ExeCoder, an LLM specifically designed for code translation, aimed at utilizing executability representations such as functional semantics, syntax structures, and variable dependencies to enhance the capabilities of LLMs in code translation. To evaluate the effectiveness of ExeCoder, we manually enhanced the widely used benchmark TransCoder-test, resulting in a benchmark called TransCoder-test-X that serves LLMs. Evaluation of TransCoder-test-X indicates that ExeCoder achieves state-of-the-art performance in code translation, surpassing existing open-source code LLMs by over 10.88% to 38.78% and over 27.44% to 42.97% on two metrics, and even outperforms the renowned closed-source LLM GPT-4o. Website: https://execoder4trans.github.io/
Large Language Model, Code Translation, Instruction Tuning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software maintenance tools
1. Introduction
Code translation aims to convert code written in one programming language into another. Translation between different programming languages can assist developers in adapting applications to new business and environments, demonstrating significant demand and value in real industrial contexts. For example, the Commonwealth Bank of Australia spent around \$750 million and five years converting its platform from COBOL to Java (Roziere et al., 2020).
@@ 1. Introduction
Figure 1. Executability Representation for Code Translation. Existing models simply copy variable name from the source code, ignoring conflicts with called built-in function, which leads to TypeError exception. The model that considers code executability learn the syntactic structure of source code and avoid call conflicts through explicit loop summation.
@@
Given the high labor costs associated with code translation, the development of automated translation tools has been extensively researched. Recent approaches (Huang et al., 2023; Roziere et al., 2021; Wen et al., 2022; Zhu et al., 2022b) utilize the paradigm of neural machine translation (NMT), aiming to learn the mapping from source language sequences to target language sequences. Despite achieving some success, their effectiveness heavily relies on the availability of cross-language parallel data (Ahmad et al., 2021b). For most languages, parallel resources are scarce or entirely absent. To overcome the limitations of NMT-based approaches, some studies (Yin et al., 2024; Macedo et al., 2024; Yang et al., 2024) have explored the use of large language models (LLMs) for code translation. These LLMs are pre-trained on a vast array of open-source code, generating code by learning the contextual semantics of the code and demonstrating excellent understanding across various programming languages. Nevertheless, existing research indicates that the correct translation rate of advanced LLMs ranges from 2.1% to 47.3%, resulting in 15 different types of execution errors (Pan et al., 2024). Leveraging the potential of LLMs for code translation poses significant challenges.
@@
In our view, the fundamental reason for this issue is the discrepancy between the pre-training tasks of existing LLMs and the requirements of code translation, as illustrated in Figure 1. Unlike natural languages, programming languages possess additional information that indicates the execution state of the code, including more complex functionalities, syntax, and variables, referred to as executability information (Jiao et al., 2023). Executability information pertains to the implementation logic of the source code, the acquisition of which relies on code analysis tools designed by programming language experts and cannot be directly inferred from the context of the source code. Code translation requires that the generated code executes correctly as intended. However, existing LLMs only learn the contextual semantics of code during pre-training, neglecting the executability information closely related to the execution state of the code, which compromises the executability of the generated code and prevents reliable automated code translation.
@@
To address these issues and harness the potential of LLMs in code translation, we propose ExeCoder in this paper. This is an LLM specifically designed for code translation, aimed at enhancing the ability of LLMs in this task by leveraging executability representations such as functional semantics, syntactic structures, and variable dependencies within the code. To extract executability knowledge from the code, ExeCoder first devises a representation strategy for executability knowledge. ExeCoder employs external code analysis tools to acquire three types of executability knowledge from the source code and meticulously designs encoding strategies to convert this knowledge into text that is easily interpretable by LLMs. To learn executability knowledge from the code, ExeCoder introduces a Progressive Executability Representation Learning strategy (PERL). The key idea is that the functional semantics, syntactic structures, and variable dependencies of the source code represent progressively refined executability information, and this staged, progressive learning aligns with the learning theories of programming experts, facilitating representation learning.
@@
To evaluate the effectiveness of ExeCoder, we conducted evaluation on the widely used code translation benchmark, TransCoder-test (Roziere et al., 2020). However, TransCoder-test can only evaluate specific implementations of the code. To address this limitation, we enhanced the TransCoder-test benchmark, resulting in a new benchmark called TransCoder-test-X, capable of evaluating the code translation capabilities of LLMs. Evaluation on TransCoder-test-X indicates that ExeCoder achieves SOTA performance in code translation tasks, surpassing existing open-source large models by over 10.88% to 38.78% and over 27.44% to 42.97% on two metrics, respectively. Notably, the ExeCoder outperforms renowned closed-source LLMs, including GPT-3.5, GPT-4, and GPT-4o, highlighting the significant role of executability representations in code translation.
@@
Figure 2. The ExeCoder’s pipeline. The ExeCoder first utilizes a text that describes the code functionality to encode functional semantic information. Subsequently, the ExeCoder uses an abstract syntax tree (AST) to encode the syntactic structure information. The AST has proven effective in encoding both lexical and syntactic information. Finally, the ExeCoder employs a data flow graph (DFG) to encode the variable dependency information. The DFG tracks the data dependency relationships between variables. With the benefit of the well-designed Progressive Executability Representation Learning (PERL) strategy, the ExeCoder can fully leverage the progressively refined code executability representation to enhance the cross-programming language understanding capabilities of LLMs.
2. Related Work
2.1. Translation of Programming Languages
@@
The translation between programming languages has been an active research field (Liu et al., 2024; Eniser et al., 2024; Khan et al., 2024; Zhu et al., 2024; Luo et al., 2024), recent work has been based on the paradigm of neural machine translation. TransCoder (Roziere et al., 2020) is the most classic unsupervised machine translation method, specifically designed with pre-training tasks to learn the semantics of source code. Other works aim to leverage code analysis tools to enhance the capabilities of code translation. SDA-Trans (Liu et al., 2023) parses the program’s syntax tree from the source code to acquire knowledge of the syntactic structure. TransCoder-IR (Szafraniec et al., 2023) parses its low-level representation, LLVM IR, from the source code and designs corresponding pre-training tasks to improve code representation. As LLMs demonstrate exceptional capabilities, their potential in code translation has also been investigated. TRANSAGENT (Yuan et al., 2024) utilizes LLMs as agents to rectify syntax and semantic errors in code translation. LASSI (Dearing et al., 2024) proposes a self-enhancement method, providing feedback during compilation to the LLM through guided debugging and refactoring prompts. Inspired by these works, ExeCoder utilizes code analysis tools to acquire knowledge related to code executability and injects this knowledge into large language models to enhance their code translation capabilities. ExeCoder has meticulously designed representation strategies for this knowledge in LLMs and proposed learning strategies for representation based on programming language learning theory to facilitate the acquisition of executability knowledge by LLMs.
2.2. Large Language Models for Code
@@
With the development of LLMs, language models tailored for code have garnered significant attention in the community. Representative LLMs include CodeLlama (Roziere et al., 2023), Qwen-Coder (Hui et al., 2024), StarCoder (Li et al., 2023), CodeT5+ (Wang et al., 2023), and Deepseek-Coder (Guo et al., 2024). These LLMs benefit from extensive pre-training on large code corpora, resulting in a strong understanding of the semantics of source code. Nevertheless, compared to state-of-the-art closed-source LLMs such as GPT-4 (Achiam et al., 2023), these models often lag in capability. To further enhance the ability of code LLMs to address specific coding issues, other works have focused on fine-tuning of pre-trained LLMs. Recent efforts include InstructCoder (Li et al., 2024), WizardCoder (Luo et al., 2023), PanGu-Coder2 (Shen et al., 2023), Magicoder (Wei et al., 2024), WaveCoder (Yu et al., 2024),. These LLMs are built upon pre-trained code LLMs and undergo post-training to further enhance model capabilities. However, these code LLMs are not specifically tailored for code translation, they merely learn from source code while neglecting the requirements for executability inherent in code translation. Unlike these works, ExeCoder encodes executability representation of source code and then employs instruction-tuning to compel LLMs to learn these executability representations, improving their code translation abilities.
3. Methodology
3.1. Overview
In this paper, we introduce ExeCoder, a large language model tailored for code translation task, which seeks to improve the capabilities of LLMs in code translation by leveraging executability representations. Executability representation refers to information indicating the execution state of the source code, which relies on code analysis tools designed by programming language experts and cannot be derived directly from the source code. The automatic translation between different programming languages necessitates that the translated code executes correctly. However, current LLMs are pre-trained only on source code, neglecting executability representations closely associated with execution states, leading to inadequate cross-programming language understanding and unreliable automated code translation. To address this problem, ExeCoder first encodes the functional semantic, syntactic structure, and variable dependency of the source code, and then employs instruction fine-tuning to compel LLMs to learn these executability representations, thereby enhancing their cross-programming language understanding and achieving more reliable automated code translation. Figure 2 illustrates the pipeline of ExeCoder.
@@ 3.1. Overview
Figure 3. Three types of Executability Representations. The first example illustrates an error related to functional semantic, where the baseline model is not informed of the function’s role in obtaining the minimum number, substituting a similar modulus symbol for the division operator. The second example highlights an error in syntactic structure, where the baseline model uses forced type conversion to convert characters to numbers, which raises a ValueError exception when non-numeric characters are included in the input. The third example presents an error regarding variable dependency, where the baseline model has not learned the transmission of variables, thereby neglecting to create a concatenation of string with itself. These errors result in a minimal edit distance but have a significant impact on the execution. Learning the executability representation can indicate the execution status, aiding in resolving these issues.
3.2. Executability Representation for Code Translation
In order to achieve more reliable automated code translation, the ExeCoder first customizes an executability knowledge representation strategy specifically for LLMs, encoding three types of executability representations of the source code, including functional semantic, syntactic structure, and variable dependency. Then, the ExeCoder constructs a specialized instruction fine-tuning dataset, XLCoST-Instruct, based on the cross-programming language alignment dataset XLCoST (Zhu et al., 2022a). XLCoST-Instruct compels LLMs to learn the executability representations of the source code, enhancing the cross-language understanding of LLMs.
@@ 3.2. Executability Representation for Code Translation
ExeCoder encodes the functional semantics of source code using a natural language description of its functionality, as shown in Figure 11. LLMs are pre-trained on a large amount of natural language data, which enables them to have a good understanding of natural language, facilitating their learning of functional semantics (Zhao et al., 2023). In fact, natural language descriptions that articulate the functional semantics of code are readily available. On one hand, a substantial amount of public code data resources is collected from open-source repositories (such as GitHub), where comments accompanying commits serve as high-quality functional semantics of code (Ahmad et al., 2021a). On the other hand, LLMs demonstrate exceptional capabilities in code summarization tasks, making it cost-effective and feasible to use external LLMs to generate natural language descriptions of source code functionality (Haldar and Hockenmaier, 2024).
@@
Syntactic-structure Representation. Compared to natural languages, programming languages have clear syntactic structures and strict grammatical rules. Merely learning the context of source code while neglecting its syntactic structure may result in translated code using incorrect syntax, leading to unexpected execution results or compilation errors, as shown in Figure 3. To address this issue, ExeCoder encodes the syntactic structure information of source code using Abstract Syntax Tree (AST) to enhance LLMs’ understanding of the syntax of different programming languages. An AST is a tree that represents the abstract syntactic structure of source code, where each subtree represents a continuous range of subword tokens, and each leaf node represents a single token; this structure has been shown to effectively encode the syntactic structure of source code (Gong et al., 2024). To construct the AST, ExeCoder utilizes the lightweight multilingual parser tree-sitter (Brunsfeld et al., 2025) to parse the source code.
@@
However, AST is structured graph data, and LLMs are pre-trained only on unstructured text, resulting in limited understanding of structured data, making it difficult to learn the syntactic structure of source code directly from the graph structure (Tang et al., 2024). To address this, inspired by (Fatemi et al., 2024), ExeCoder further encodes the AST into unstructured text that is easier for LLMs to understand, serving as a representation of the syntactic structure, as shown in Figure 12. Specifically, ExeCoder first removes the content of non-leaf nodes in the AST, retaining only the tokens on the leaf nodes. This strategy aims to eliminate redundant code segments and reduce noise in the representation of syntactic structure. Next, ExeCoder treats the processed AST as a graph and assigns a numerical index to each node. Finally, ExeCoder represents the node information and edge information of the graph using natural language. For the node information, ExeCoder describes the token content of each node. For the edge information, ExeCoder describes the neighboring nodes of each node. Following this strategy, the syntactic structure information of the source code is ultimately encoded into a segment of natural language. Existing work (Fatemi et al., 2024) demonstrates that this representation strategy significantly enhances LLMs’ understanding of graph structure information.
@@
Variable-dependency Representation. In different programming languages, the same variable often has different semantics due to variations in programmer preferences and naming conventions (Cheng et al., 2024). The diversity of variable semantics may lead to incorrect variable dependency relationships when solely learning the semantics of source code, further resulting in anomalous or unexpected execution outcomes, as shown in Figure 3. To address this issue, ExeCoder encodes the variable dependency information of the source code using Data Flow Graph (DFG) to enhance LLMs’ understanding of variable dependencies in code. A DFG is a graph that describes the dependencies and interactions between variables, where each node represents a variable and each edge indicates the source of these variables. To construct the DFG, ExeCoder utilizes the lightweight multilingual parser tree-sitter (Brunsfeld et al., 2025) to parse the source code.
@@
3.3. XLCoST-Instruct
To enable LLMs to learn the executability representation and enhance their cross-language understanding capabilities, ExeCoder developed a specially designed instruction fine-tuning dataset XLCoST-Instruct for code translation. Details of XLCoST-Instruct are in Appendix B and C.
3.4. Progressive Executability Representation Learning
ExeCoder aims to utilize executability representations such as functional semantics, syntactic structures, and variable dependencies in code to enhance the capabilities of LLMs in code translation tasks. However, the acquisition of executability representations relies on code analysis tools, which, unlike source code, are domain-specific high-level knowledge that is difficult to comprehend directly, facing two significant challenges in the learning process. First, how can we learn both low-level code semantics and high-level executability representations simultaneously? Second, how can we design learning strategies to enhance the understanding of high-level executability representations? To address these issues, ExeCoder proposes a Progressive Executability Representation Learning (PERL) to leverage progressively refined executability representations of code to enhance LLMs’ cross-program language understanding capabilities.
Specifically, to address the first issue, PERL incorporates executability representations as auxiliary knowledge during instruction fine-tuning, as illustrated in Appendix C. PERL concatenates the source code with its corresponding executability representation and then prompts LLMs to translate code based on the respective executability representations. This strategy has been shown to effectively enable LLMs to learn high-level knowledge (Yin et al., 2023).
To address the second issue, PERL has designed a phased fine-tuning strategy. Research (Robins et al., 2003) indicates that the optimal learning of programming skills is progressive, beginning with an understanding of the high-level intentions of a program, followed by the implementation of low-level code structures and behaviors. Inspired by this, PERL has designed a phased fine-tuning strategy where each fine-tuning stage independently learns different executability representations. The key idea is that the functional semantics, syntactic structures, and variable dependencies of source code are executability representations that are refined progressively. Phased progressive learning aligns with the learning theories of programming experts, facilitating representational learning. Specifically, ExeCoder divides the instruction fine-tuning process into four stages: source code, functional semantics, syntactic structures, and variable dependencies, with each stage fine-tuning using only one type of data. Once the fine-tuning process of one stage converges, the next stage of fine-tuning is initiated.
4. Experimental Evaluation
Table 1. Execution-based evaluation of ExeCoder and baseline models on TransCoder-test-X.

### v3

@@
Abstract
Code translation is a crucial activity in the software development and maintenance process, and researchers have recently begun to focus on using pre-trained large language models (LLMs) for code translation. However, existing LLMs only learn the contextual semantics of code during pre-training, neglecting executability information closely related to the execution state of the code, which results in unguaranteed code executability and unreliable automated code translation. To address this issue, we propose ExeCoder, an LLM specifically designed for code translation, aimed at utilizing executability representations such as functional semantics, syntax structures, and variable dependencies to enhance the capabilities of LLMs in code translation. To evaluate the effectiveness of ExeCoder, we manually enhanced the widely used benchmark TransCoder-test, resulting in a benchmark called TransCoder-test-X that serves LLMs. Evaluation of TransCoder-test-X indicates that ExeCoder achieves state-of-the-art performance in code translation, surpassing existing open-source code LLMs by over 10.88% to 38.78% and over 27.44% to 42.97% on two metrics, and even outperforms the renowned closed-source LLM GPT-4o. Code is available at https://aka.ms/execoder
ExeCoder: Empowering Large Language Models with Executability Representation for Code Translation
Minghua Hethanks: These authors contributed equally to this work. thanks: This work was done during the internship at Microsoft.1, Yue Chen11footnotemark: 122footnotemark: 21, Fangkai Yangthanks: Corresponding author.2, Pu Zhao2, Wenjie Yin3, Yu Kang2, Qingwei Lin2, Saravan Rajmohan2, Dongmei Zhang2 1Peking University, 2Microsoft, 3KTH Royal Institute of Technology hemh2120@stu.pku.edu.cn, fangkai.yang@microsoft.com
1 Introduction
Code translation aims to convert code written in one programming language into another. Translation between different languages can assist developers in adapting applications to new business and environments, demonstrating significant demand and value in real industrial contexts. For example, the Commonwealth Bank of Australia spent around \$750 million and five years converting its platform from COBOL to Java Roziere et al. (2020).
@@ 1 Introduction
Figure 1: Executability Representation for Code Translation. Existing models simply copy variable name from the source code, ignoring conflicts with called built-in function, which leads to TypeError exception. The model that considers code executability learn the syntactic structure of source code and avoid call conflicts through explicit loop summation.
@@
Given the high labor costs associated with code translation, the development of automated translation tools has been extensively researched. Recent approaches Huang et al. (2023); Roziere et al. (2021); Wen et al. (2022); Zhu et al. (2022b) utilize the paradigm of neural machine translation (NMT), aiming to learn the mapping from source language sequences to target language sequences. Despite achieving some success, their effectiveness heavily relies on the availability of cross-language parallel data Ahmad et al. (2021a). For most languages, parallel resources are scarce or entirely absent. To overcome the limitations of NMT-based approaches, some studies Yin et al. (2024); Lu et al. (2025); Mei et al. (2025); Macedo et al. (2024); Yang et al. (2024) have explored the use of large language models (LLMs) for code translation. These LLMs are pre-trained on a vast array of open-source code, generating code by learning the contextual semantics of the code and demonstrating excellent understanding across various programming languages. Nevertheless, existing research indicates that the correct translation rate of advanced LLMs ranges from 2.1% to 47.3%, resulting in 15 different types of execution errors Pan et al. (2024). Leveraging the potential of LLMs for code translation poses significant challenges.
@@
In our view, the fundamental reason for this issue is the discrepancy between the pre-training tasks of existing LLMs and the requirements of code translation, as illustrated in Figure 1. Unlike natural languages, programming languages possess additional information that indicates the execution state of the code, including more complex functionalities, syntax, and variables, referred to as executability information Jiao et al. (2023). Executability information pertains to the implementation logic of the source code, the acquisition of which relies on code analysis tools designed by programming language experts and cannot be directly inferred from the context of the source code. Code translation requires that the generated code executes correctly as intended. However, existing LLMs only learn the contextual semantics of code during pre-training, neglecting the executability information closely related to the execution state of the code, which compromises the executability of the generated code and prevents reliable automated code translation.
@@
To address these issues and harness the potential of LLMs in code translation, we propose ExeCoder. This is an LLM specifically designed for code translation, aimed at enhancing the ability of LLMs by leveraging executability representations such as functional semantics, syntactic structures, and variable dependencies within the code. To extract executability knowledge from the code, ExeCoder first devises a representation strategy for executability knowledge. ExeCoder employs external code analysis tools to acquire three types of executability knowledge from the source code and meticulously designs encoding strategies to convert this knowledge into text that is easily interpretable by LLMs. To learn executability knowledge from the code, ExeCoder introduces a Progressive Executability Representation Learning strategy (PERL). The key idea is that the functional semantics, syntactic structures, and variable dependencies of the source code represent progressively refined executability information, and this staged, progressive learning aligns with the learning theories of programming experts, facilitating representation learning.
@@
To evaluate the effectiveness of ExeCoder, we conducted evaluation on the widely used code translation benchmark, TransCoder-test Roziere et al. (2020). However, TransCoder-test can only evaluate specific implementations of the code. To address this limitation, we enhanced the TransCoder-test benchmark, resulting in a new benchmark called TransCoder-test-X, capable of evaluating the code translation capabilities of LLMs. Evaluation on TransCoder-test-X indicates that ExeCoder achieves SOTA performance in code translation tasks, surpassing existing open-source large models by over 10.88% to 38.78% and over 27.44% to 42.97% on two metrics, respectively. Notably, the ExeCoder outperforms renowned closed-source LLMs, including GPT-3.5, GPT-4, and GPT-4o, highlighting the significant role of executability representations in code translation.
@@
Figure 2: The ExeCoder’s pipeline. The ExeCoder first utilizes a text that describes the code functionality to encode functional semantic information. Subsequently, the ExeCoder uses an abstract syntax tree (AST) to encode the syntactic structure information. Finally, the ExeCoder employs a data flow graph (DFG) to encode the variable dependency information. With the benefit of the well-designed Progressive Executability Representation Learning (PERL) strategy, the ExeCoder can fully leverage the executability representation to enhance the cross-language understanding capabilities of LLMs.
2 Related Work
2.1 Translation of Programming Languages
@@
The translation between programming languages has been an active research field Liu et al. (2024); Eniser et al. (2024); Khan et al. (2024); Zhu et al. (2024); Luo et al. (2024), recent work has been based on the paradigm of neural machine translation. TransCoder Roziere et al. (2020) is the most classic unsupervised machine translation method, specifically designed with pre-training tasks to learn the semantics of source code. Other works aim to leverage code analysis tools to enhance the capabilities of code translation. SDA-Trans Liu et al. (2023) parses the program’s syntax tree from the source code to acquire knowledge of the syntactic structure. TransCoder-IR Szafraniec et al. (2023) parses its low-level representation, LLVM IR, from the source code and designs corresponding pre-training tasks to improve code representation. As LLMs demonstrate exceptional capabilities, their potential in code translation has also been investigated. TRANSAGENT Yuan et al. (2024) utilizes LLMs as agents to rectify syntax and semantic errors in code translation. LASSI Dearing et al. (2024) proposes a self-enhancement method, providing feedback during compilation to the LLM through guided debugging and refactoring prompts. Inspired by these works, ExeCoder utilizes code analysis tools to acquire knowledge related to code executability and injects this knowledge into LLMs to enhance their code translation capability.
2.2 Large Language Models for Code
@@
With the development of LLMs, language models tailored for code have garnered significant attention in the community. Representative LLMs include CodeLlama Roziere et al. (2023), Qwen-Coder Hui et al. (2024), StarCoder Li et al. (2023), CodeT5+ Wang et al. (2023), and Deepseek-Coder Guo et al. (2024). These LLMs benefit from extensive pre-training on large code corpora, resulting in a strong understanding of the semantics of source code. Nevertheless, compared to state-of-the-art closed-source LLMs such as GPT-4 Achiam et al. (2023), these models often lag in capability. To further enhance the ability of code LLMs to address specific coding issues, other works have focused on fine-tuning of pre-trained LLMs. Recent efforts include InstructCoder Li et al. (2024), WizardCoder Luo et al. (2023), PanGu-Coder2 Shen et al. (2023), Magicoder Wei et al. (2024), WaveCoder Yu et al. (2024),. These LLMs are built upon pre-trained code LLMs and undergo post-training to further enhance model capabilities. However, these code LLMs are not specifically tailored for code translation, they merely learn from source code while neglecting the requirements for executability inherent in code translation. Unlike these works, ExeCoder encodes executability representation of source code and then employs instruction-tuning to compel LLMs to learn these executability representations, improving their code translation abilities.
3 Methodology
@@
Figure 3: Three types of Executability Representations. The first example illustrates an error related to functional semantic, where the baseline model is not informed of the function’s role in obtaining the minimum number, substituting a similar modulus symbol for the division operator. The second example highlights an error in syntactic structure, where the baseline model uses forced type conversion to convert characters to numbers, which raises a ValueError exception when non-numeric characters are included in the input. The third example presents an error regarding variable dependency, where the baseline model has not learned the transmission of variables, thereby neglecting to create a concatenation of string with itself. These errors result in a minimal edit distance but have a significant impact on the execution. Learning the executability representation can indicate the execution status, aiding in resolving these issues.
3.1 Overview
In this paper, we introduce ExeCoder, a LLM tailored for code translation task, which seeks to improve the capabilities of LLMs in code translation by leveraging executability representations. Executability representation refers to information indicating the execution state of the source code, which relies on code analysis tools designed by programming language experts and cannot be derived directly from the source code. The automatic translation between different programming languages necessitates that the translated code executes correctly. However, current LLMs are pre-trained only on source code, neglecting executability representations closely associated with execution states, leading to inadequate cross-programming language understanding and unreliable automated code translation. To address this problem, ExeCoder first encodes the executability representations of the source code, and then employs instruction fine-tuning to compel LLMs to learn these executability representations, thereby enhancing their cross-programming language understanding and achieving more reliable automated code translation. Figure 2 illustrates the pipeline of ExeCoder.
3.2 Executability Representation for Code Translation
In order to achieve more reliable code translation, the ExeCoder first customizes an executability knowledge representation strategy specifically for LLMs, encoding three types of executability representations, including functional semantic, syntactic structure, and variable dependency. Then, the ExeCoder constructs a specialized instruction fine-tuning dataset, XLCoST-Instruct, based on the cross-programming language alignment dataset XLCoST Zhu et al. (2022a). XLCoST-Instruct compels LLMs to learn the executability representations of the source code, enhancing the cross-language understanding of LLMs.
@@ 3.2 Executability Representation for Code Translation
ExeCoder encodes the functional semantics of source code using a natural language description of its functionality, as shown in Figure 11. LLMs are pre-trained on a large amount of natural language data, which enables them to have a good understanding of natural language, facilitating their learning of functional semantics Zhao et al. (2023). In fact, natural language descriptions that articulate the functional semantics of code are readily available. A substantial amount of public code data resources is collected from open-source repositories (such as GitHub), where comments accompanying commits serve as high-quality functional semantics of code Ahmad et al. (2021b).
@@
Syntactic-structure Representation. Compared to natural languages, programming languages have clear syntactic structures and strict grammatical rules. Merely learning the context of source code while neglecting its syntactic structure may result in translated code using incorrect syntax, leading to unexpected execution results or compilation errors, as shown in Figure 3. To address this issue, ExeCoder encodes the syntactic structure information of source code using Abstract Syntax Tree (AST) to enhance LLMs’ understanding of the syntax of different programming languages. An AST is a tree that represents the abstract syntactic structure of source code, where each subtree represents a continuous range of subword tokens, and each leaf node represents a single token; this structure has been shown to effectively encode the syntactic structure of source code Gong et al. (2024). To construct the AST, ExeCoder utilizes the lightweight multilingual parser tree-sitter Brunsfeld et al. (2025) to parse the source code.
@@
However, AST is structured graph data, and LLMs are pre-trained only on unstructured text, resulting in limited understanding of graph structure Tang et al. (2024), making it difficult to learn the syntactic structure of source code. To address this, inspired by Fatemi et al. (2024), ExeCoder further encodes the AST into unstructured text that is easier for LLMs to understand, serving as a representation of the syntactic structure, as shown in Figure 12. ExeCoder simplifies ASTs by keeping only leaf node tokens. It then indexes the AST as a graph and describes nodes and edges in natural language, enhancing LLM understanding of code structure Fatemi et al. (2024). Further details are provided in Appendix D.2.
@@
Variable-dependency Representation. In different programming languages, the same variable often has different semantics due to variations in programmer preferences and naming conventions Cheng et al. (2024). The diversity of variable semantics may lead to incorrect variable dependency relationships when solely learning the semantics of source code, further resulting in anomalous or unexpected execution outcomes, as shown in Figure 3. To address this issue, ExeCoder encodes the variable dependency information of the source code using Data Flow Graph (DFG) to enhance LLMs’ understanding of variable dependencies in code. A DFG is a graph that describes the dependencies and interactions between variables, where each node represents a variable and each edge indicates the source of these variables. To construct the DFG, ExeCoder utilizes the lightweight multilingual parser tree-sitter Brunsfeld et al. (2025) to parse the source code.
@@
3.3 XLCoST-Instruct
To enable LLMs to learn the executability representation and enhance their cross-language understanding capabilities, ExeCoder developed a specially designed instruction fine-tuning dataset XLCoST-Instruct for code translation. Details of XLCoST-Instruct are in Appendix E and F.
3.4 Progressive Executability Representation Learning
ExeCoder aims to utilize executability representations such as functional semantics, syntactic structures, and variable dependencies in code to enhance the capabilities of LLMs in code translation tasks. However, the acquisition of executability representations relies on code analysis tools, which, unlike source code, are domain-specific high-level knowledge that is difficult to comprehend directly, facing two significant challenges in the learning process. First, how can we learn both low-level code semantics and high-level executability representations simultaneously? Second, how can we design learning strategies to enhance the understanding of high-level executability representations? To address these issues, ExeCoder proposes a Progressive Executability Representation Learning (PERL) to leverage progressively refined executability representations of code to enhance LLMs’ cross-language understanding capabilities.
Specifically, to address the first issue, PERL incorporates executability representations as auxiliary knowledge during instruction fine-tuning, as illustrated in Appendix F. PERL concatenates the source code with its corresponding executability representation and then prompts LLMs to translate code based on the respective executability representations. This strategy has been shown to effectively enable LLMs to learn high-level knowledge Yin et al. (2023).
To address the second issue, PERL has designed a phased fine-tuning strategy. Research Robins et al. (2003) indicates that the optimal learning of programming skills is progressive, beginning with an understanding of the high-level intentions of a program, followed by the implementation of low-level code structures. Inspired by this, PERL has designed a phased fine-tuning strategy where each fine-tuning stage independently learns different executability representations. The key idea is that the functional semantics, syntactic structures, and variable dependencies of source code are executability representations that are refined progressively. Phased progressive learning aligns with the learning theories of programming experts, facilitating representational learning. Specifically, ExeCoder divides the instruction fine-tuning process into four stages: source code, functional semantics, syntactic structures, and variable dependencies, with each stage fine-tuning using only one type of data. Once the fine-tuning process of one stage converges, the next stage of fine-tuning is initiated.
4 Experimental Evaluation